# Intertwined Nematicity, Multiferroicity, and Nonlinear Hall Effect in Rhombohedral Pentalayer Graphene

### v1

@@
Intertwined Nematicity, Multiferroicity, and Nonlinear Hall Effect in Rhombohedral Pentalayer Graphene
@@
(March 13, 2025)
Electronic nematicity-the spontaneous reduction of rotational symmetry-has been widely studied in strongly correlated quantum materials, yet its interplay with other symmetry-breaking phases remains an outstanding experimental challenge to unravel, particularly in van der Waals heterostructures. Here, using angle-resolved transport measurements, we reveal a unique intertwinement of nematicity, orbital multiferroicity, and the nonlinear Hall effect in rhombohedral pentalayer graphene. This coupling enables electric-field-driven switching of the nematic principal axis and the nonlinear Hall polar vector, coinciding with the butterfly-shaped hysteresis of the multiferroic order. We identify two sharply defined nematic phase transitions: the first coincides with the onset of orbital ferromagnetism, while the second, at lower temperatures, aligns with the emergence of ferroelectricity. Our findings point to a nematic-driven instability that intertwines rotational, time-reversal, and inversion symmetry breaking. This is consistent with spontaneous momentum polarization, proposed to occur under the influence of the Coulomb-driven flocking effect. Together, our observations establish nematicity as a fundamental organizing principle in the landscape of intertwined electronic orders in strongly correlated two-dimensional systems.
Figure 1: Rhombohedral-stacking pentalayer graphene and angle-resolved transport measurement. (a) Longitudinal resistance $R_{\parallel}$ as a function of carrier density $n$ and displacement field $D$ in a dual-encapsulated, dual-gated rhombohedral pentalayer graphene device. The inset shows schematic diagram of the heterostructure. (b) Schematic of the angle-resolved transport measurement setup used to extract $R_{\parallel}$ (left) and $R_{\perp}$ (right) in a disk-shaped sample. (c) Angular dependence of $R_{\parallel}$ and $R_{\perp}$ inside the triangular multiferroic region, measured at $T=40$ mK and $B=0$ T. (d) $R_{0}$ and (e) $R_{H}$, extracted from the angular dependence using Eq. 1-2, measured as $B$ is swept back and forth.
@@
Figure 3: Nematic phase transitions. (a-b) Temperature dependence of (a) the magnitude of the anomalous Hall coefficient $|R_{H}|$ and (b) the strength of transport anisotropy $\Delta R/R_{0}$. Upper right inset shows a zoomed in plot of $\Delta R/R_{0}$ around $T_{nem}^{1}$, where the dashed curve denote the best fit using the Curie-Weiss law. (c-e) Angular dependence of $R_{\parallel}$ and $R_{\perp}$ measured (c) in temperature regime i at $T=10$ K, (d) in temperature regime ii at $T=2.8$ K, and (e) in temperature regime iii at $T=0.4$ K. (f-h) Anomalous Hall coeffecient $R_{H}$ measured with sweeping $D$ back and forth (f)) in temperature regime i at $T=5$ K, (g) in temperature regime ii at $T=3.5$ K, and (h) in temperature regime iii at $T=0.04$ K.
@@
In strongly interacting quantum materials, the interplay of correlations, broken symmetry, and topology gives rise to a complex landscape of intertwined electronic orders at low temperature  Fradkin et al. (2015); Keimer and Moore (2017); Fernandes et al. (2019); Nuckolls and Yazdani (2024); Wang et al. (2023). Among these, nematicity has been shown to play a pivotal role in shaping the low-temperature phase diagram  Kivelson et al. (1998); Fradkin et al. (2010); Oganesyan et al. (2001); Fernandes et al. (2014); Böhmer et al. (2022); Cao et al. (2020); Zhang et al. (2024). Extensive research has revealed its deep connection to competing phases. In iron-based superconductors, for example, nematicity frequently coexists with magnetic and superconducting transitions, underscoring a fundamental yet still poorly understood connection between these emergent orders  Chu et al. (2012); Fernandes et al. (2014); Böhmer et al. (2022). Despite its ubiquity across material platforms, the precise nature of nematicity remains elusive in most quantum materials, especially low-dimensional systems, primarily due to experimental challenges in isolating and characterizing this interaction-driven order.
This raises several fundamental questions: What constitutes definitive evidence of Coulomb-driven nematicity in transport measurement, distinct from lattice distortions  Fernandes et al. (2014); Böhmer et al. (2022); Ren et al. (2021); Calvera et al. (2024)? Can nematicity be controlled and switched, similar to other Coulomb-driven instabilities, such as ferromagnetism and ferroelectricity  Samajdar et al. (2021); Naritsuka et al. (2023); Klein et al. (2020)? What is the precise relationship between nematicity, other symmetry-breaking orders, and topological phases  Samkharadze et al. (2016); Sun et al. (2009)? Resolving these challenges requires experimental strategies capable of identifying intrinsic nematic order using transport method—a crucial step toward unraveling the mechanisms governing electronic symmetry breaking and enabling new approaches to manipulate emergent quantum phenomena.
In this work, we investigate the low-temperature phase diagram of rhombohedral-stacked pentalayer graphene (rPG), a system where inherently flat energy bands and nontrivial topology lead to a striking interplay between ferromagnetism and ferroelectricity, culminating in the emergence of multiferroicity  Han et al. (2023, 2024); Dong et al. (2024); Lu et al. (2024). Using angle-resolved transport measurement in both the linear and nonlinear regimes  Wu et al. (2017, 2020); Vafek (2023); Chichinadze et al. (2024); Zhang et al. (2024), we establish a previously unrecognized connection between nematicity, the nonlinear Hall effect  Sodemann and Fu (2015); Ma et al. (2019); Kang et al. (2019); Chichinadze et al. (2024), and multiferroicity  Han et al. (2023). This coupling manifests in electric-field-driven switching and hysteretic behaviors in both nematicity and the nonlinear Hall effect, demonstrating unprecedented control over electronic symmetry breaking. Additionally, our results uncover a pair of thermodynamic transitions, driven by nematic instabilities, that intertwine rotational, time-reversal, and inversion symmetry breaking, imposing stringent constraints on the underlying electronic order in rhombohedral graphene.
@@
We begin by investigating the transport properties of rPG in the linear regime, where an ohmic response is probed using a small d.c. current bias of $I_{dc}=5$ nA. Fig. 1a plots the longitudinal resistance $R_{\parallel}$ across the low-temperature phase space, parameterized by charge carrier density ($n$) and displacement field ($D$). The resulting $n-D$ map exhibits high symmetry around $D=0$, consistent with the absence of substrate-induced sublattice symmetry breaking, indicative of misalignment between rPG and hBN. Notably, a triangular region emerges near $D=0$ on the hole-doping side of the charge-neutrality (CNP), hosting an exotic multiferroic order at low temperature  Han et al. (2023). The remainder of this report focuses on unraveling the complex intertwinement of competing orders within this region.
@@
To perform angle-resolved transport measurements  Chichinadze et al. (2024); Zhang et al. (2024), the rPG sample is patterned in a “sunflower” geometry, featuring a disk-shaped area with eight evenly spaced electrical contacts, as illustrated in Fig. 1b. Transport measurements are conducted by applying a current bias between two contacts, designated as source (S) and drain (D), with their relative orientation defining the azimuth direction of current flow ($\phi$). By measuring the voltage response between contacts $A$ and $B$, we extract the longitudinal ($R_{\parallel}$) and transverse ($R_{\perp}$) resistance.
@@
Here, $\Delta R$ represents the oscillation amplitude, $R_{0}$ is the average value of $R_{\parallel}$$(\phi)$, $R_{H}$ is the average value of $R_{\perp}$$(\phi)$, and $\alpha$ denotes the principal axis orientation, defined by the current direction of the highest conductivity. According to theoretical discussion  Vafek (2023), Eq. 1-2 fully determines the conductivity matrix describing transport response in the linear regime: the ratio between maximum and minimum conductivity, qualified by $\Delta R/R_{0}$, offers a direct measure of anisotropy strength. Furthermore, a non-zero $R_{H}$ indicates the presence of antisymmetric off-diagonal terms in the conductivity matrix, corresponding to an anomalous Hall coefficient associated with orbital ferromagnetism.
@@
In addition to its coupling to magnetism, the anomalous Hall effect exhibits electric-field-driven switching, a hallmark of multiferroic order Han et al. (2023). Figures 2a–c show that a large anomalous Hall coefficient ($R_{H}$) emerges in the low-density portion of the triangular region, with red and blue regions indicating opposite Hall responses. A sign reversal in $R_{H}$ occurs across $D=0$, suggesting that the Hall response is controlled by the displacement field. Notably, $R_{H}$ is highly sensitive to the direction of $D$ sweeping: at a fixed point in the $n$$D$ phase space, reversing the $D$ sweep direction induces a corresponding sign reversal in $R_{H}$, giving rise to an electric-field-driven hysteresis in the anomalous Hall effect.
@@
To further elucidate this switching, Fig. 2h–i compare the angular dependence of transport measured at $D=25$ mV/nm, correspondng to the red and blue open circles in Fig. 2d-g. When $D$ is swept backward (from positive to negative), the transport response exhibits a large angular oscillation, indicating strong anisotropy, accompanied by a positive $R_{H}$. In contrast, as $D$ is swept forward (from negative to positive), a reduced $\Delta R/R_{0}$ signals a suppression of transport anisotropy, while the principal axis ($\alpha$) undergoes a significant rotation. Notably, at negative $D$, strong (weak) anisotropy occurs with forward (backward) $D$-sweeping (see Fig. XX), opposite to the behavior observed in Fig. 2h–i, indicating a strong coupling between electronic anisotropy and the ferroelectric switching.
This hysteretic switching in transport anisotropy, driven by an electric-field, demonstrates unprecedented control over the emergence of rotational symmetry breaking. Such control is fundamentally distinct from the effects of strain and distortion in the crystal lattice, suggesting that the observed transport anisotropy is primarily of electronic origin.
Furthermore, the electric-field-driven switching also underscores the intrinsic connection between broken rotational symmetry, valley polarization, and layer polarization. Due to the abrupt changes in both $\Delta R/R_{0}$ and $\alpha$, the transport response measured along certain azimuth angles does not fully capture the electric-field-driven transitions. For instance, along $\phi=22.5^{\circ}$, $R_{\perp}$ does not exhibit the characteristic butterfly-shaped hysteresis (see Fig. M3). In this context, angle-resolved transport serves as an essential tool for unraveling the unique intertwinement between nematicity and multiferroicity.
To further disentangle these intertwined orders, we analyze their temperature dependence. Fig. 3a-b plot the temperature dependence of $|R_{H}|$ and $\Delta R/R_{0}$, extracted from angle-resolved transport measurements (see Method Section for measurement details). This analysis reveals three distinct temperature regimes, separated by two transitions, denoted as $T_{\text{nem}^{1}}$ and $T_{\text{nem}^{2}}$, respectively.
@@
The temperature dependence of $\Delta R/R_{0}$ just above $T_{\text{nem}^{1}}$ is well approximated by a Curie-Weiss curve, diverging at $T_{0}=2.1$ K (inset in Fig. 3b). This behavior closely resembles nematic phase transitions observed in other quantum materials, such as Iron-based superconductors  Chu et al. (2012); Fernandes et al. (2014); Böhmer et al. (2022). This temperature dependence provides strong evidence for a thermodynamic transition driven by a nematic instability, marking a previously unobserved phenomenon in 2D van der Waals materials.
Furthermore, the simultaneous emergence of nematicity and orbital ferromagnetism at $T_{\text{nem}}^{1}$ suggests a common underlying mechanism that intertwines rotational and time-reversal symmetry breaking. This behavior aligns with theoretical predictions of Coulomb-driven instabilities in momentum space, where a collective flocking effect induces spontaneous momentum condensation  Dong et al. (2023); Jung et al. (2015); Huang et al. (2023); Mandal and Fernandes (2023); Parra-Martinez et al. (2025).
@@
The nonlinear transport response is extracted by measuring the voltage response at the second harmonic frequency of an AC current bias  Ma et al. (2019); Kang et al. (2019); He et al. (2022); Zhang et al. (2024); Chichinadze et al. (2024). To ensure measurements remain within the quadratic response regime, we apply a small AC current of $10-15$ nA (see Methods section for details). Fig. 4a-b present the angular dependence of the nonlinear response inside the multiferroic regime, revealing hallmark signatures of the nonlinear Hall effect. Both the parallel ($V_{\parallel}^{2\omega}$) and perpendicular ($V_{\perp}^{2\omega}$) channels of the nonlinear response exhibit a one-fold angular oscillation, with $V_{\perp}^{2\omega}$ displaying a significantly larger oscillation amplitude. This behavior is in excellent agreement with the predicted angular response of the nonlinear Hall effect  Chichinadze et al. (2024). Most remarkably, our findings indicate that the nonlinear Hall component plays a dominating role in the conductivity tensor describing transport response in the nonlinear regime.
@@
Taken together, the observed correlation between nematicity, the nonlinear Hall effect, and orbital ferromagnetism provides unambiguous evidence for a Coulomb-driven instability that intertwines rotational, time-reversal, and inversion symmetry breaking. This phenomenon is in excellent agreement with the theoretical framework of spontaneous momentum polarization.
@@
Having examined the nature of $T_{\text{nem}}^{1}$, we now turn to the low-temperature transition at $T_{\text{nem}}^{2}$, which coincides with the emergence of ferroelectricity. The fact that the anomalous Hall effect and electric-field-driven switching emerge at distinct temperatures suggests that ferromagnetism and ferroelectricity originate from different mechanisms, despite occupying the same region of the $n$$D$ phase space at low temperatures.
@@
Furthermore, the coincidence of electric-field-driven hysteresis with the sharp onset in $\Delta R/R_{0}$ provides compelling evidence that ferroelectric order is intrinsically anisotropic, a crucial property not accounted for in previous theoretical models of multiferroicity in this system Das and Huang (2024).
The interplay between ferroelectricity and rotational symmetry breaking is further reflected in the tunability of the nonlinear Hall effect at $T<T_{nem}^{2}$. Figures 4g–h plot the angular dependence of $V_{\perp}^{2\omega}$, measured at a fixed carrier density ($n$) and displacement field ($D$), revealing a prominent rotation of the nonlinear Hall vector ($\mathcal{C}$) as $D$ is swept in opposite directions (Fig. 4f). This $D$-induced hysteresis not only establishes an unprecedented degree of control over the nonlinear Hall effect, but also underscores an intrinsic link with nematicity and multiferroicity.
@@
Lastly, we investigate the fermiology of the multiferroic state by analyzing the Shubnikov–de Haas (SdH) quantum oscillations. As shown in Fig. 5a, quantum oscillations exhibit a “fish-net” pattern across a wide carrier density range. We perform a Fourier transform of $R_{\parallel}$ as a function of $1/B$ to identify the the Luttinger volume and degeneracy of the Fermi sea responsible for the “fish-net” pattern (Fig. 5b).
@@
Figure 5c marks the center frequencies of the most prominent peaks in the Fourier spectrum, revealing distinct density regimes characterized by different Fermi surface topologies. At low carrier density ($n=-0.5\times 10^{12}$ cm-2), the two dominant frequencies, $f_{1}$ and $f_{2}$, satisfy $f_{1}-f_{2}=1$, suggesting two possible Fermi surface topologies: (i) a quarter metal phase characterized by an annular Fermi sea or (ii) an exotic half-metallic state, where an electron pocket coexists with a hole pocket, each featuring opposite valley polarizations (inset of Fig. 5c).
[…]
@@
Notably, angle-resolved transport measurements reveal no anomalous Hall effect in the low- and high-density regimes (Fig. 2b-c). This is inconsistent with a quarter metal phase at low-density, suggesting that the Fermi surface instead consists of coexisting electron and hole pockets with opposite valley polarizations, where the overall Chern number remains close to zero. The universal presence of electron and hole pockets is in line with prior theoretical predictions Das and Huang (2024). In the presence of an out-of-plane electric field, the electron and hole pockets polarize to opposite graphene layers, creating a distinct charge distribution that likely serves as the direct link between nematicity and ferroelectricity.
@@
During the transition between these high- and low-density regimes, the spectral weight of the Fourier transform shifts from $f_{1}$ and $f_{2}$ to $f_{3}$ and $f_{4}$, signaling the onset of a partially isospin-polarized (PIP) phase  Zhou et al. (2021); Holleis et al. (2025). Notably, the emergence of multiferroicity at $B=0$ coincides with the density range of the PIP phase, suggesting that Fermi surface reconstruction plays a pivotal role in driving momentum-space instabilities. These instabilities, in turn, are responsible for the simultaneous breaking of time-reversal, rotational, and inversion symmetries.
@@
According to prior theoretical discussions, spontaneous momentum polarization could arise from the instability of an annular Fermi sea  Huang et al. (2023); Jung et al. (2015), or from exchange-driven condensation between discrete Fermi pockets  Dong et al. (2023); Parra-Martinez et al. (2025). These mechanisms provide a compelling framework for understanding the interplay between Fermi surface topology, isospin physics, and emergent ferroic orders in rhombohedral graphene.
Overall, through angle-resolved transport measurements, we uncover the simultaneous emergence of nematicity, multiferroicity, and the nonlinear Hall effect, establishing an unprecedented scenario where rotational, time-reversal, and inversion symmetries are broken in tandem. These findings provide strong experimental support for the Coulomb-driven phenomenon of spontaneous momentum-space condensation  Huang et al. (2023); Jung et al. (2015); Dong et al. (2023); Parra-Martinez et al. (2025), which has been theoretically discussed in connection with the exotic chiral superconductivity.

### v2

# Coulomb-driven Momentum Space Condensation in Rhombohedral Hexalayer Graphene

@@
Coulomb-driven Momentum Space Condensation in Rhombohedral Hexalayer Graphene
@@
(August 8, 2025)
In systems of strongly correlated electrons, the emergence of spontaneous symmetry breaking serves as a defining hallmark of the underlying many-body physics. Here, we report the discovery of an unprecedented emerging order in rhombohedral hexalayer graphene, defined by the intertwinement of rotational, time-reversal, and inversion symmetry breaking. The link between these broken symmetries is uncovered through angle-resolved transport measurements. The spontaneous breaking of rotational symmetry is identified by the Curie-Weiss behavior in the temperature-dependent onset of transport anisotropy. Most remarkably, the nematic transition, defined by the cusp of transport anisotropy, coincides with the emergence of the anomalous Hall effect in both the linear and nonlinear regimes. Our findings provide unambiguous experimental evidence of spontaneous condensation in momentum space, a phenomenon previously proposed to arise from a Coulomb-driven flocking effect. Together, these observations establish electronic anisotropy as a fundamental organizing principle in the landscape of intertwined electronic orders in strongly correlated two-dimensional systems.
Figure 1: Rhombohedral-stacking hexalayer graphene and angle-resolved transport measurement. (a) Longitudinal resistance $R_{\parallel}$ as a function of carrier density $n$ and displacement field $D$ in a dual-encapsulated, dual-gated rhombohedral hexalayer graphene device. The inset shows schematic diagram of the heterostructure. (b) Schematic of the angle-resolved transport measurement setup used to extract $R_{\parallel}$ (left) and $R_{\perp}$ (right) in a disk-shaped sample. (c) Angular dependence of $R_{\parallel}$ and $R_{\perp}$ inside the triangular multiferroic region, measured at $T=40$ mK and $B=0$ T. (d) $R_{0}$ and (e) $R_{H}$, extracted from the angular dependence using Eq. 1-2, measured as $B$ is swept back and forth.
@@
Figure 3: Nematic phase transitions. (a-b) Temperature dependence of (a) the magnitude of the anomalous Hall coefficient $|R_{H}|$ and (b) the strength of transport anisotropy $\Delta R/R_{0}$. Upper right inset shows a zoomed in plot of $\Delta R/R_{0}$ around $T_{nem}^{1}$, where the dashed curve denote the best fit using the Curie-Weiss law. (c-e) Angular dependence of $R_{\parallel}$ and $R_{\perp}$ measured (c) in temperature regime i at $T=10$ K, (d) in temperature regime $i$ at $T=2.8$ K, and (e) in temperature regime iii at $T=0.4$ K. (f-h) Anomalous Hall coeffecient $R_{H}$ measured with sweeping $D$ back and forth (f)) in temperature regime i at $T=5$ K, (g) in temperature regime $i$ at $T=3.5$ K, and (h) in temperature regime iii at $T=0.04$ K.
@@
Rhombohedral multilayer graphene systems have recently emerged as a compelling platform for exploring strongly correlated quantum phenomena  Zhou et al. (2021); Han et al. (2024a); Chen et al. (2020); Han et al. (2023); Lu et al. (2024); Choi et al. (2025); Liu et al. (2024); Sha et al. (2024); Han et al. (2024b). While the energy bands remain dispersive at high carrier densities, the application of a perpendicular electric field ($D$) dramatically enhances the density of states near the band edges, producing a broad range of interaction-driven phases. The resulting low-temperature phase space hosts a diverse spectrum of emergent orders, many of which exhibit spontaneous time-reversal symmetry breaking—including half- and quarter-metallic states  Zhou et al. (2021); Han et al. (2024a); Chen et al. (2020), multiferroicity  Han et al. (2023), and various quantum Hall effect states at zero external magnetic field  Lu et al. (2024); Choi et al. (2025); Liu et al. (2024); Sha et al. (2024); Han et al. (2024b).
Beyond time-reversal symmetry breaking, another critical component of the underlying electronic order has been proposed in several theoretical works  Jung et al. (2015); Huang et al. (2023); Dong et al. (2023). These studies suggest that the quarter-metal phase is accompanied by spontaneous breaking of $C_{3}$ rotational symmetry. Driven by a Coulomb-induced flocking instability, charge carriers spontaneously condense into a single corner of the trigonally-warped Fermi surface, acquiring a finite net momentum. This mechanism links broken time-reversal and rotational symmetries. The resulting electronic ground state, described by momentum polarization, may play a foundational role in shaping the broader phase diagram of rhombohedral multilayer graphene.
Despite its theoretical appeal, direct experimental identification of a momentum-polarized ground state has remained elusive, largely due to the intrinsic difficulty of detecting spontaneous rotational symmetry breaking using conventional probes  Zhang et al. (2024); Lin et al. (2023). Efforts to characterize electronic orders with such broken symmetries raise several fundamental questions: What constitutes definitive evidence of spontaneous rotational symmetry breaking, distinct from lattice-induced distortions  Fernandes et al. (2014); Böhmer et al. (2022); Ren et al. (2021); Calvera et al. (2024)? Can this broken symmetry be externally controlled and switched, in analogy to other Coulomb-driven instabilities such as ferromagnetism Samajdar et al. (2021); Naritsuka et al. (2023); Klein et al. (2020)? And finally, what are the distinct transport signatures of an electronic state that simultaneously breaks time-reversal and rotational symmetries  Huang et al. (2023)?
Motivated by these questions, in this work we investigate the low-temperature phase diagram of rhombohedral-stacked hexalayer graphene (RHG). Using angle-resolved transport measurements in both the linear and nonlinear regimes Wu et al. (2017, 2020); Vafek (2023); Chichinadze et al. (2024); Zhang et al. (2024), we uncover a unique intertwinement between transport anisotropy, the nonlinear Hall effect Sodemann and Fu (2015); Ma et al. (2019); Kang et al. (2019); Chichinadze et al. (2024), and orbital multiferroicity Han et al. (2023). Together, these intertwined orders provide unambiguous experimental support for a momentum-polarized state that simultaneously breaks time-reversal, rotational, and inversion symmetries, offering new insight into the nature of Coulomb-driven symmetry breaking in rhombohedral multilayer graphene.
@@
We begin by investigating the transport properties of RHG in the linear regime, where an ohmic response is probed using a small d.c. current bias of $I_{dc}=5$ nA. Fig. 1a plots the longitudinal resistance $R_{\parallel}$ across the low-temperature phase space, parameterized by charge carrier density ($n$) and displacement field ($D$). The resulting nDn-D map exhibits high symmetry around $D=0$, consistent with the absence of substrate-induced sublattice symmetry breaking, indicative of misalignment between RHG and hBN. Notably, a triangular region emerges near $D=0$ on the hole-doping side of the charge-neutrality (CNP), hosting an exotic multiferroic order at low temperature  Han et al. (2023). The remainder of this report focuses on unraveling the nature of spontaneous symmetry breaking within this region using angle-resolved transport.
@@
To perform angle-resolved transport measurements  Chichinadze et al. (2024); Zhang et al. (2024), the RHG sample is patterned in a “sunflower” geometry, featuring a disk-shaped area with eight evenly spaced electrical contacts, as illustrated in Fig. 1b. Transport measurements are conducted by applying a current bias between two contacts, designated as source (S) and drain (D), with their relative orientation defining the azimuth direction of current flow ($\phi$). By measuring the voltage response between contacts $A$ and $B$, we extract the longitudinal ($R_{\parallel}$) and transverse ($R_{\perp}$) resistance.
@@
Here, $\Delta R$ represents the oscillation amplitude, $R_{0}$ is the average value of $R_{\parallel}$$(\phi)$, $R_{H}$ is the average value of $R_{\perp}$$(\phi)$, and $\alpha$ denotes the principal axis orientation, defined by the current direction of the highest conductivity. According to theoretical discussion  Vafek (2023), Eq. 1-2 fully determines the conductivity matrix describing transport response in the linear regime: the ratio between the oscillation amplitude and average value, $\Delta R/R_{0}$, offers a direct measure of anisotropy strength. Furthermore, a non-zero $R_{H}$ indicates the presence of antisymmetric off-diagonal terms in the conductivity matrix, corresponding to an anomalous Hall coefficient associated with orbital ferromagnetism.
@@
In addition to its coupling to magnetism, the anomalous Hall effect exhibits electric-field-driven switching, a hallmark of multiferroic order Han et al. (2023). Figures 2a–c show that a large anomalous Hall coefficient ($R_{H}$) emerges in the low-density portion of the triangular region, with red and blue indicating opposite Hall responses. A sign reversal in $R_{H}$ occurs across $D=0$, suggesting that the Hall response is controlled by the displacement field. Notably, $R_{H}$ is highly sensitive to the direction of $D$ sweeping: at a fixed point in the nnDD phase space, reversing the $D$ sweep direction induces a corresponding sign reversal in $R_{H}$, giving rise to an electric-field-driven hysteresis in the anomalous Hall effect.
@@
To further elucidate this switching, Fig. 2h–i compare the angular dependence of transport measured at $D=25$ mV/nm, corresponding to the red and blue open circles in Fig. 2d-g. When $D$ is swept backward (from positive to negative), the transport response exhibits a large angular oscillation, indicating strong anisotropy, accompanied by a positive $R_{H}$. In contrast, as $D$ is swept forward (from negative to positive), a reduced $\Delta R/R_{0}$ signals a suppression of transport anisotropy, while the principal axis ($\alpha$) undergoes a significant rotation. Notably, at negative $D$, strong (weak) anisotropy occurs with forward (backward) $D$-sweeping (see Fig. M1), opposite to the behavior observed in Fig. 2h–i, indicating a strong coupling between electronic anisotropy and the ferroelectric switching.
The hysteretic switching in transport anisotropy, driven by an electric-field, demonstrates an unprecedented level of control over rotational symmetry breaking. Unlike strain- or distortion-induced effects in the crystal lattice, this control arises purely from electronic interactions, indicating that the observed anisotropy is fundamentally of electronic origin.
Because both $\Delta R/R_{0}$ and $\alpha$ undergo abrupt changes across the transition, transport responses measured along a fixed azimuth angle may not fully capture the electric-field-driven symmetry breaking. For example, at $\phi=22.5^{\circ}$, $R_{\perp}$ does not display the characteristic butterfly-shaped hysteresis (see Fig. M3). In this context, angle-resolved transport measurements are essential for uncovering the symmetry and nature of the underlying electronic order in phases with strong transport anisotropy.
Next, we analyze the temperature-dependent onset of rotational and time-reversal symmetry breaking. Fig. 3a-b plot the temperature dependence of $|R_{H}|$ and $\Delta R/R_{0}$, extracted from angle-resolved transport measurements. This analysis reveals three distinct temperature regimes, separated by two transitions, denoted as $T_{\text{nem}^{1}}$ and $T_{\text{nem}^{2}}$, respectively.
@@
The temperature dependence of $\Delta R/R_{0}$ just above $T_{\text{nem}^{1}}$ is well approximated by a Curie-Weiss curve (black dashed line in the inset of Fig. 3b), diverging at $T_{0}=2.1$ K. This behavior closely resembles nematic phase transitions observed in other quantum materials, such as Iron-based superconductors  Chu et al. (2012); Fernandes et al. (2014); Böhmer et al. (2022). This temperature dependence provides strong support for the presence of a nematic instability, marking a previously unobserved phenomenon in 2D van der Waals materials.
Furthermore, the emergence of the anomalous Hall effect at $T_{\text{nem}}^{1}$ suggests a common underlying mechanism that intertwines rotational and time-reversal symmetry breaking. This behavior aligns with theoretical predictions of Coulomb-driven instabilities in momentum space, where a collective flocking effect induces spontaneous momentum condensation  Dong et al. (2023); Jung et al. (2015); Huang et al. (2023); Mandal and Fernandes (2023); Parra-Martinez et al. (2025).
@@
The nonlinear transport response is extracted by measuring the voltage response at the second harmonic frequency of an AC current bias  Ma et al. (2019); Kang et al. (2019); He et al. (2022); Zhang et al. (2024); Chichinadze et al. (2024). To ensure measurements remain within the quadratic response regime, we apply a small AC current of 101510-15 nA (see Methods section for details). Fig. 4a-b present the angular dependence of the nonlinear response inside the multiferroic regime, revealing hallmark signatures of the nonlinear Hall effect. Both the parallel ($V_{\parallel}^{2\omega}$) and perpendicular ($V_{\perp}^{2\omega}$) channels of the nonlinear response exhibit a one-fold angular oscillation, with $V_{\perp}^{2\omega}$ displaying a significantly larger oscillation amplitude. This behavior is in excellent agreement with the predicted angular response of the nonlinear Hall effect  Chichinadze et al. (2024). Most remarkably, our findings indicate that the nonlinear Hall component plays a dominating role in the conductivity tensor describing transport response in the nonlinear regime.
@@
Taken together, the observed correlation between transport anisotropy, the nonlinear Hall effect, and orbital ferromagnetism provides unambiguous evidence for a Coulomb-driven instability that intertwines rotational, time-reversal, and inversion symmetry breaking. This phenomenon is in excellent agreement with the theoretical framework of spontaneous momentum polarization.
@@
Further decreasing temperature unlocks another transition at $T_{\text{nem}}^{2}$, which coincides with the emergence of ferroelectricity. The fact that the anomalous Hall effect and electric-field–driven switching emerge at distinct temperatures suggests that ferromagnetism and ferroelectricity originate from different mechanisms, despite occupying the same region of the nnDD phase space at low temperatures.
@@
At the same time, the coincidence of electric-field–driven hysteresis with the sharp onset in $\Delta R/R_{0}$ provides compelling evidence that ferroelectric order is intrinsically anisotropic, a crucial property not accounted for in previous theoretical models of multiferroicity in this system Das and Huang (2024).
The presence of ferroelectricity at $T<T_{\text{nem}}^{2}$ also gives rise to electric-field–driven tunability in the nonlinear Hall effect. Figures 4g–h plot the angular dependence of $V_{\perp}^{2\omega}$, measured at a fixed carrier density ($n$) and displacement field ($D$). As $D$ is swept in opposite directions, angle-resolved measurement of the nonlinear response uncovers a rotation of $90^{\circ}$ in the nonlinear Hall vector ($\mathcal{C}$) (Fig. 4f). Such a prominent rotation suggests that momentum polarization is decoupled from sample details, such as lattice-induced distortion, allowing it to be freely controlled using the interplay with ferroelectricity.
@@
Lastly, we investigate the fermiology of the multiferroic state by analyzing the Shubnikov–de Haas (SdH) quantum oscillations. As shown in Fig. 5a, quantum oscillations exhibit a “fish-net” pattern across a wide carrier density range. We perform a Fourier transform of $R_{\parallel}$ as a function of $1/B$ to identify the the Luttinger volume and degeneracy of the underlying Fermi sea (Fig. 5b).
@@
Figure 5c marks the center frequencies of the most prominent peaks in the Fourier spectrum, revealing distinct density regimes characterized by different Fermi surface topologies. At low carrier density ($n=-0.5\times 10^{12}$ cm-2), the two dominant frequencies, $f_{1}$ and $f_{2}$, satisfy $f_{1}-f_{2}=1$, suggesting two possible Fermi surface topologies: (i) a quarter metal characterized by an annular Fermi sea or (ii) an exotic half-metal, where an electron pocket coexists with a hole pocket, each featuring opposite valley polarizations (inset of Fig. 5c).
At higher densities ($2.5\times 10^{12}$ $^{-2}<n<-1.0\times 10^{12} cm-$2), the sum rule $f_{3}-f_{4}=1/2$ suggests either (i) a half-metal phase with an annular Fermi sea or (ii) the coexistence of two electron and two hole pockets, indicative of a more complex isospin texture.
@@
Since angle-resolved transport measurements reveal no anomalous Hall effect in the low- and high-density regimes (Fig. 2b-c), the Fermi surface likely consists of coexisting electron and hole pockets with opposite valley polarizations, where the overall Chern number remains close to zero. This is in line with prior theoretical predictions Das and Huang (2024). In the presence of an out-of-plane electric field, the electron and hole pockets polarize to opposite graphene layers, creating a distinct charge distribution that likely serves as the direct link between transport anisotropy and ferroelectricity.
@@
As the multiferroic order emerges during the transition between these high- and low-density regimes, the spectral weight of the Fourier transform shifts from $f_{1}$ and $f_{2}$ to $f_{3}$ and $f_{4}$, signaling the onset of a partially isospin-polarized (PIP) phase Zhou et al. (2021); Holleis et al. (2025). This suggests that Fermi surface reconstruction plays a pivotal role in enabling momentum-space instabilities, which, in turn, are responsible for the simultaneous breaking of time-reversal, rotational, and inversion symmetries.
@@
Overall, our angle-resolved transport measurements in RHG establish an unprecedented scenario where rotational, time-reversal, and inversion symmetries are broken in tandem. These findings provide strong experimental support for the Coulomb-driven phenomenon of spontaneous momentum-space condensation  Huang et al. (2023); Jung et al. (2015); Dong et al. (2023); Parra-Martinez et al. (2025). According to prior theoretical discussions, momentum polarization could arise from the instability of an annular Fermi sea Huang et al. (2023); Jung et al. (2015), or from exchange-driven condensation between discrete Fermi pockets Dong et al. (2023); Parra-Martinez et al. (2025). These mechanisms provide a compelling framework for understanding the interplay between Fermi surface topology, isospin physics, and emergent ferroic orders in rhombohedral graphene systems.